OpenStudy (anonymous):

Simplify the expression. 34 + 12 ÷ 3 - (1-9) a.)75 b.)39 c.)93 d.)21

OpenStudy (anonymous):

did you state your question correctly

OpenStudy (anonymous):

yes . . .

OpenStudy (anonymous):

34+12/3 -(1-9)?

OpenStudy (anonymous):

34+4+8=46

OpenStudy (anonymous):

Simplify the expression. 34 + 12 ÷ 3 - (1-9)

OpenStudy (anonymous):

Simplify. 8 + 3 [3-(-1)6]

OpenStudy (anonymous):

Simplify the expression. 34 + 12 ÷ 3 - (1-9)

OpenStudy (anonymous):

ientered it correctLy !

OpenStudy (anonymous):

hisbarby... Please see the attached file as it describes it better. The right answer would be 46 as dichalao mentioned it on top.

OpenStudy (anonymous):

hisbarbie, you have to be careful when you enter your expressions.

OpenStudy (anonymous):

the 4 is a exponent sorry !

OpenStudy (anonymous):

See, lol!

OpenStudy (anonymous):

93

OpenStudy (anonymous):

\[3^4+(12/3)-(1-9)=81+4-(-8)=81+4+8=93\]

OpenStudy (anonymous):

thanks so much !

OpenStudy (anonymous):

np

OpenStudy (anonymous):

good luck